# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v5

@@ 6.2.1 Object detection and instance segmentation
We present a comparative analysis of SOTA, a self-supervised approaches for downstream tasks such as object detection and instance segmentation (as shown in Table 5. For an unbiased evaluation, we compare techniques having ViT-B [11] as the backbone architecture. The methods are trained over COCO [110] dataset considering Mask R-CNN [129] as the task head for object detection and instance segmentation and evaluated using average precision $AP^{box}$ and $AP^{mask}$, respectively. The results in Table 5 demonstrate that TEC [119], dBOT [120], and MSG-MAE [116] outperform other SOTA SSL methods over object detection and obtain $AP^{box}$ scores of 52.3, 52.7, and 52.3 respectively. On the instance segmentation, MSG-MAE [116], TEC [119], and dBOT [120] show significant promising performance compared to other SSL methods and achieve 48.8, 47.2, and 45.7 $AP^{mask}$ scores, respectively.
@@ 6.2.2 Semantic Segmentation Comparison
In Table 5, we present a performance comparison of SSL methods for semantic segmentation, where UpperNet [130] is trained with ViT-B [11] over ADE20K [131] dataset. From Table 5, it is notable that BootMAE [122], Diet-III [123], Diet-III+AugMask [124], MSG-MAE [116], TEC [119] and dBOT [120] exhibit above 49.0% mIoU. We also observe that top performing Diet-III+AugMask [124], TEC [119], Diet-III [123], and MSG-MAE [116], demonstate the 50.2%, 49.9%, 49.7%, and 49.7%, respectively.
@@
DINO [61] ViT-B/16 99.1 91.7 72.6 78.6 98.8 93.0
MSN [132] ViT-B/16 99.0 90.5 72.1 78.1 - -
iBOT [117] ViT-B/16 99.2 92.2 74.6 79.6 98.9 94.3
MAE [69] ViT-B/16 - - 75.4 80.5 - -
dBOT [120] ViT-B/16 99.3 91.3 77.9 81.0 98.2 93.7
MoCo-v3 [93] ViT-B/16 98.9 90.5 - 97.7 -
MixedAE [127] ViT-B/16 97.9 85.9 - - 97.1 88.8
Diet-III [123] ViT-B/16 99.3 92.5 73.6 78.0 98.6 93.4
@@ 6.2.3 Transfer Classification Comparison
To further study the generalizability of the SSL methods, we compare their transfer learning capabilities in terms of accuracy. To do so, the pre-trained methods are fine-tuned over small datasets such as CIFAR-10 [111], CIFAR-100 [111], iNaturalist18 [133] (iNa18), iNaturalist19 [133] (iNa19), Flowers [134], and Cars [135]. The results in Table 6 Deit-III [123] demonstrate outstanding performance over state-of-the-art SSL methods and achieve 99.3 and 92.5 accuracy over CIFAR-10 and CIFAR-100 datasets, accordingly. We also note that dBOT [120] illustrates better performance over CIFAR-10, iNa18, and iNa19 datasets, and show accuracy of 99.3, 77.9, and 81.0, respectively. Similarly, iBOT [117] exhibits the best performance over Flowers and Cars datasets while obtaining 98.9 and 94.3 accuracy, respectively.
@@ 6.3 Robustness to Perturbations
We also present a comparative comparison of SSL methods over various robustness datasets including natural adversarial examples [136] (IN-A), objects in different styles and textures (IN-R [137]), controls in rotation, background, and viewpoints (ObjNet [138]), and SI-scores [139] (SI-size, SI-loc, and SI-rot). The results in Table 7 show that Diet-III+AugMask [124] has better robustness across various robustness metrics.
Table 7: Robustness comparison over various robustness benchmarks using ViT-B as a backbone network. It is noticeable that Diet-III+AugMask [124] demonstrates better robustness performance. The results in bold indicate better performance.
@@
Diet-III [123] 83.8 73.4 88.2 36.8 54.1 35.7 58.0 42.7 41.5
Diet-III+AugMask [124] 84.2 74.0 88.6 41.9 54.4 37.2 59.0 44.8 43.3
MFM [126] 32.7 48.6
MAE [69] 31.5 48.3
@@ 7.1.1 Traditional Augmentation
Augmentation using geometric transformation is being used for various image processing applications. By employing this transformation, the original positions of the image pixels are relocated to new positions without modifying intensity values. It modifies the training data by incorporating variations in viewpoint, non-rigid deformations, and scaling for better model learning on real world problems. The geometric transformations including the translation, rotation, reflection, shearing, and scaling are being commonly used for data augmentation. It is categorized into affine and non-affine transformations. The former one utilized linear mapping for adjusting the geometric disorder in the structure of the image. Whereas the later involved complex non-linear mapping functions [140, 141]. The projective and perspective transformations are fall into the non-affine geometric transformations. These are effectively being used in computer vision and medical imaging tasks [142, 143]. Similarly, the deformation transformations are non-affine which are used to simulate at higher degree of freedom such as Lens and non-rigid body deformations [144, 145, 146]. Photometric transformation [147, 148] is also fall under the traditional augmentation category and used for various computer vision tasks. It considers the camera and shooting artifacts such as optical noise, motion blur, image color artifacts degradations. These simulations are being used to mitigate the effects of data scarcity during model training.
@@ 7.1.2 Advanced Augmentation Strategies
The advanced augmentation strategies are used to solve complex computer vision problems and gained a significant attention of researchers. These augmentation operations are learned from the given data automatically. Several CNN based models are proposed in literature to perform various advanced transformations for data augmentation [149, 150, 151]. These deep models have several cascading layers to learn the transformations, consequently obtained a better trained and generalized model. Jagerberg et al. (2015) [149] proposed spatial transformer network (STN) data augmentation approach which is used for model training. Based on STN data augmentation model, several improved models are built which offer promising results [152, 153, 154]. Although the advanced models are computationally expensive compared to the traditional methods. However, the accuracy and diversity of data augmentation of these models are very high. They are effective to solve various data scarcity related computer vision and SSL problems.
@@ 7.1.3 Challenges of Data Augmentation
  • Designing of augmentation strategies: Generally, traditional methods for data augmentation techniques includes flipping, random cropping, and color jittering. It is not a simple task to design a strategy for effective data augmentation in SSL. Exploring novel and appropriate augmentations which exploit certain characteristics of ViTs is an open challenge [155].
  • Robustness to complex variations: Usually, real-world images having complex variations such as in lighting, occlusion, viewpoint, etc. It is a major challenge to validate that during SSL and data augmentation process such complex variations are truly accounted [156].
  • Maintaining of consistency and discriminability: Typically, SSL techniques use produced augmented visions of the same image and capture useful patterns to support consistency that is small changes in the input will reflect in the learned representations. In order to maintain discriminative information, it is challenging to discover the sufficient balance between produced augmented visions consistent for learning [157].
  • Scale Sensitivity: ViTs work excellent on large-scale datasets. The data augmentation methods well-behaved on data scarcity problems and it might not be effective on larger-scale datasets. Development of augmentation methods, which are capable to apply on larger-scale datasets is a challenge [11].
  • Computational Efficiency: Commonly, ViTs are computationally intensive, and incorporation of data through augmentation techniques would make problem more intensified. Exploring adequate data augmentation methods which are computationally efficient and offering valuable benefits to SSL is a practical challenge [158].
  • Robust against adversarial attacks: Data augmentation methods usually are not vulnerable against adversarial attacks. It is open challenge to develop efficient data augmentation approaches which improve developed model robustness against various forms of adversarial attacks [159].
@@ 7.2 Incorporating Spatial Information
Incorporating spatial information into SSL-based ViTs is an active area of research nowadays. Spatial details are essential for grasping the structural relationships within images and can improve the model’s ability to capture meaningful features. One recent trend incorporating spatial information is utilizing positional embeddings using sine/cosine and axial position embeddings [160]. Spatial relationships can also be better captured using grid masking [161]. The attention mechanisms can also be used to capture the spatial dependencies, e.g., axial attention [162] which can help the model to focus on local and global spatial features. However, incorporating spatial information might be challenging due to increase the computational complexity of transformers. Further, it complicates when dealing with multi-modal data or tasks that require understanding spatial-temporal relationships. Therefore, spatial attention is one of the components of SSL and ViTs that can be encountered when solving a complex problem.
@@ 7.3 Improving Sample Efficiency
The performance of sample efficiency can be improved by following some of the good practices. For instance, bigger training datasets can be further divided into subsets to capture fine details of input data. If the training dataset is already small, then labeled data can be combined with unlabeled data to perform SSL learning. Regularization methods such as dropout, weight decay, and layer normalization can also improve sample efficiency. Regularization helps to prevent overfitting and improve the model’s generalization. It is recommended to employ the transfer learning concept from larger, pre-trained models for training of smaller ViTs to mimic the behaviour of the larger models [163]. By following this practice, smaller model’s performance can be improved while having limited dataset. Fine-tuning strategies such as layer freezing, gradual unfreezing, and learning rate schedules can also affect the sample efficiency.
@@ 7.4 Interpretability and Explainability
Interpretability and explainability are critical for the utilization and veracity of the models. To understand the inner working mechanism of the model while solving real world problems, the following steps might be considered while designing the ViTs models. For interpretability, the important aspects include feature visualization, activation analysis, and attribution methods can be considered during model development [164, 165, 166] . While, for explainability of ViTs models, the attention mechanism, attention maps, feature attribution, concept activation, and shapely values need to be accounted [167, 168, 169].

### v6

@@ 6.2.1 Object detection and instance segmentation
We present a comparative analysis of SOTA, a self-supervised approaches for downstream tasks such as object detection and instance segmentation (as shown in Table 5. For an unbiased evaluation, we compare techniques having ViT-B (Dosovitskiy et al. 2020) as the backbone architecture. The methods are trained over COCO (Lin et al. 2014) dataset considering Mask R-CNN (Cai and Vasconcelos 2019) as the task head for object detection and instance segmentation and evaluated using average precision $AP^{box}$ and $AP^{mask}$, respectively. The results in Table 5 demonstrate that TEC (Gao et al. 2022), dBOT (Liu et al. 2022), and MSG-MAE (Tukra et al. 2023) outperform other SOTA SSL methods over object detection and obtain $AP^{box}$ scores of 52.3, 52.7, and 52.3 respectively. On the instance segmentation, MSG-MAE (Tukra et al. 2023), TEC (Gao et al. 2022), and dBOT (Liu et al. 2022) show significant promising performance compared to other SSL methods and achieve 48.8, 47.2, and 45.7 $AP^{mask}$ scores, respectively.
@@ 6.2.2 Semantic Segmentation Comparison
In Table 5, we present a performance comparison of SSL methods for semantic segmentation, where UpperNet (Xiao et al. 2018) is trained with ViT-B (Dosovitskiy et al. 2020) over ADE20K (Zhou et al. 2017) dataset. From Table 5, it is notable that BootMAE (Dong et al. 2022), Diet-III (Touvron et al. 2022), Diet-III+AugMask (Heo et al. 2023), MSG-MAE (Tukra et al. 2023), TEC (Gao et al. 2022) and dBOT (Liu et al. 2022) exhibit above 49.0% mIoU. We also observe that top performing Diet-III+AugMask (Heo et al. 2023), TEC (Gao et al. 2022), Diet-III (Touvron et al. 2022), and MSG-MAE (Tukra et al. 2023), demonstate the 50.2%, 49.9%, 49.7%, and 49.7%, respectively.
@@
DINO (Caron et al. 2021) ViT-B/16 99.1 91.7 72.6 78.6 98.8 93.0
MSN (Assran et al. 2022) ViT-B/16 99.0 90.5 72.1 78.1 - -
iBOT (Zhou et al. 2021) ViT-B/16 99.2 92.2 74.6 79.6 98.9 94.3
MAE (He et al. 2022) ViT-B/16 - - 75.4 80.5 - -
dBOT (Liu et al. 2022) ViT-B/16 99.3 91.3 77.9 81.0 98.2 93.7
MoCo-v3 (Chen* et al. 2021) ViT-B/16 98.9 90.5 - 97.7 -
MixedAE (Chen et al. 2023) ViT-B/16 97.9 85.9 - - 97.1 88.8
Diet-III (Touvron et al. 2022) ViT-B/16 99.3 92.5 73.6 78.0 98.6 93.4
@@ 6.2.3 Transfer Classification Comparison
To further study the generalizability of the SSL methods, we compare their transfer learning capabilities in terms of accuracy. To do so, the pre-trained methods are fine-tuned over small datasets such as CIFAR-10 (Krizhevsky et al. 2009), CIFAR-100 (Krizhevsky et al. 2009), iNaturalist18 (Van Horn et al. 2018) (iNa18), iNaturalist19 (Van Horn et al. 2018) (iNa19), Flowers (Nilsback and Zisserman 2008), and Cars (Krause et al. 2013). The results in Table 6 Deit-III (Touvron et al. 2022) demonstrate outstanding performance over state-of-the-art SSL methods and achieve 99.3 and 92.5 accuracy over CIFAR-10 and CIFAR-100 datasets, accordingly. We also note that dBOT (Liu et al. 2022) illustrates better performance over CIFAR-10, iNa18, and iNa19 datasets, and show accuracy of 99.3, 77.9, and 81.0, respectively. Similarly, iBOT (Zhou et al. 2021) exhibits the best performance over Flowers and Cars datasets while obtaining 98.9 and 94.3 accuracy, respectively.
@@ 6.3 Robustness to Perturbations
We also present a comparative comparison of SSL methods over various robustness datasets including natural adversarial examples (Hendrycks et al. 2021a) (IN-A), objects in different styles and textures (IN-R (Hendrycks et al. 2021b)), controls in rotation, background, and viewpoints (ObjNet (Barbu et al. 2019)), and SI-scores (Djolonga et al. 2021) (SI-size, SI-loc, and SI-rot). The results in Table 7 show that Diet-III+AugMask (Heo et al. 2023) has better robustness across various robustness metrics.
Table 7: Robustness comparison over various robustness benchmarks using ViT-B as a backbone network. It is noticeable that Diet-III+AugMask (Heo et al. 2023) demonstrates better robustness performance. The results in bold indicate better performance.
@@
Diet-III (Touvron et al. 2022) 83.8 73.4 88.2 36.8 54.1 35.7 58.0 42.7 41.5
Diet-III+AugMask (Heo et al. 2023) 84.2 74.0 88.6 41.9 54.4 37.2 59.0 44.8 43.3
MFM (Xie et al. 2022) 32.7 48.6
MAE (He et al. 2022) 31.5 48.3
@@ 7.1.1 Traditional Augmentation
Augmentation using geometric transformation is being used for various image processing applications. By employing this transformation, the original positions of the image pixels are relocated to new positions without modifying intensity values. It modifies the training data by incorporating variations in viewpoint, non-rigid deformations, and scaling for better model learning on real world problems. The geometric transformations including the translation, rotation, reflection, shearing, and scaling are being commonly used for data augmentation. It is categorized into affine and non-affine transformations. The former one utilized linear mapping for adjusting the geometric disorder in the structure of the image. Whereas the later involved complex non-linear mapping functions (Gallier 2011, Struik 2011). The projective and perspective transformations are fall into the non-affine geometric transformations. These are effectively being used in computer vision and medical imaging tasks (Wang et al. 2019, Franke et al. 2021). Similarly, the deformation transformations are non-affine which are used to simulate at higher degree of freedom such as Lens and non-rigid body deformations (Milletari et al. 2016, Simard et al. 2003, Ronneberger et al. 2015). Photometric transformation (Xu et al. 2023, Mumuni and Mumuni 2022) is also fall under the traditional augmentation category and used for various computer vision tasks. It considers the camera and shooting artifacts such as optical noise, motion blur, image color artifacts degradations. These simulations are being used to mitigate the effects of data scarcity during model training.
@@ 7.1.2 Advanced Augmentation Strategies
The advanced augmentation strategies are used to solve complex computer vision problems and gained a significant attention of researchers. These augmentation operations are learned from the given data automatically. Several CNN based models are proposed in literature to perform various advanced transformations for data augmentation (Jaderberg et al. 2015, Karargyris 2015, Tarasiuk and Pryczek 2016). These deep models have several cascading layers to learn the transformations, consequently obtained a better trained and generalized model. Jagerberg et al. (2015) (Jaderberg et al. 2015) proposed spatial transformer network (STN) data augmentation approach which is used for model training. Based on STN data augmentation model, several improved models are built which offer promising results (Mounsaveng et al. 2021, Luo et al. 2020, Vu and Huang 2017). Although the advanced models are computationally expensive compared to the traditional methods. However, the accuracy and diversity of data augmentation of these models are very high. They are effective to solve various data scarcity related computer vision and SSL problems.
@@ 7.1.3 Challenges of Data Augmentation
  • Designing of augmentation strategies: Generally, traditional methods for data augmentation techniques includes flipping, random cropping, and color jittering. It is not a simple task to design a strategy for effective data augmentation in SSL. Exploring novel and appropriate augmentations which exploit certain characteristics of ViTs is an open challenge (Cubuk et al. 2020).
  • Robustness to complex variations: Usually, real-world images having complex variations such as in lighting, occlusion, viewpoint, etc. It is a major challenge to validate that during SSL and data augmentation process such complex variations are truly accounted (Geirhos et al. 2018).
  • Maintaining of consistency and discriminability: Typically, SSL techniques use produced augmented visions of the same image and capture useful patterns to support consistency that is small changes in the input will reflect in the learned representations. In order to maintain discriminative information, it is challenging to discover the sufficient balance between produced augmented visions consistent for learning (Tian et al. 2020).
  • Scale Sensitivity: ViTs work excellent on large-scale datasets. The data augmentation methods well-behaved on data scarcity problems and it might not be effective on larger-scale datasets. Development of augmentation methods, which are capable to apply on larger-scale datasets is a challenge (Dosovitskiy et al. 2020).
  • Computational Efficiency: Commonly, ViTs are computationally intensive, and incorporation of data through augmentation techniques would make problem more intensified. Exploring adequate data augmentation methods which are computationally efficient and offering valuable benefits to SSL is a practical challenge (Touvron et al. 2019).
  • Robust against adversarial attacks: Data augmentation methods usually are not vulnerable against adversarial attacks. It is open challenge to develop efficient data augmentation approaches which improve developed model robustness against various forms of adversarial attacks (Hassan et al. 2023).
@@ 7.2 Incorporating Spatial Information
Incorporating spatial information into SSL-based ViTs is an active area of research nowadays. Spatial details are essential for grasping the structural relationships within images and can improve the model’s ability to capture meaningful features. One recent trend incorporating spatial information is utilizing positional embeddings using sine/cosine and axial position embeddings (Chu et al. 2021). Spatial relationships can also be better captured using grid masking (Li et al. 2021). The attention mechanisms can also be used to capture the spatial dependencies, e.g., axial attention (ZHENG2023105587 2023) which can help the model to focus on local and global spatial features. However, incorporating spatial information might be challenging due to increase the computational complexity of transformers. Further, it complicates when dealing with multi-modal data or tasks that require understanding spatial-temporal relationships. Therefore, spatial attention is one of the components of SSL and ViTs that can be encountered when solving a complex problem.
@@ 7.3 Improving Sample Efficiency
The performance of sample efficiency can be improved by following some of the good practices. For instance, bigger training datasets can be further divided into subsets to capture fine details of input data. If the training dataset is already small, then labeled data can be combined with unlabeled data to perform SSL learning. Regularization methods such as dropout, weight decay, and layer normalization can also improve sample efficiency. Regularization helps to prevent overfitting and improve the model’s generalization. It is recommended to employ the transfer learning concept from larger, pre-trained models for training of smaller ViTs to mimic the behaviour of the larger models (Wang et al. 2022). By following this practice, smaller model’s performance can be improved while having limited dataset. Fine-tuning strategies such as layer freezing, gradual unfreezing, and learning rate schedules can also affect the sample efficiency.
@@ 7.4 Interpretability and Explainability
Interpretability and explainability are critical for the utilization and veracity of the models. To understand the inner working mechanism of the model while solving real world problems, the following steps might be considered while designing the ViTs models. For interpretability, the important aspects include feature visualization, activation analysis, and attribution methods can be considered during model development (Zeiler and Fergus 2014, Zhu et al. 2017, Selvaraju et al. 2017) . While, for explainability of ViTs models, the attention mechanism, attention maps, feature attribution, concept activation, and shapely values need to be accounted (Koh and Liang 2017, Guidotti et al. 2018, Lundberg and Lee 2017).
@@ 9 Conclusion
Statements and Declarations
Funding
This research did not receive external funding.
Competing Interests
The authors declare no conflict of interest.